If (a,a^2)falls inside the angle made by the line y=x/2,x>0 and y=3x,x>0 then a belongs to:

Grade:11

1 Answers

Arun
6 years ago
y = x/2 ; m1 = tan -1(1/2)
y = 3x ; m2 = tan -1(3)
 
Since the point (a,a2) lies between the angle auntanded by lines L1 and L2 at origin qill lie between m1 and m2.
Hence
m = a^2 -0/ a- 0 = a
 
tan -1(1/2) -1(3)